The Faber Family has 3 children. Bill's age plus Mike's age add to 20 years. Mike's agee plus Sally's age add to 29 years. Bill's age plus Sally's age add to 23 years. The difference in age between the old child and young child is what?
Answer by ankor@dixie-net.com(22740)

The Faber Family has 3 children.
:
Write an equation for each statement:
:
Bill's age plus Mike's age add to 20 years.
b + m = 20
:
Mike's age plus Sally's age add to 29 years.
m + s = 29
:
Bill's age plus Sally's age add to 23 years.
b + s = 23
b = (-s + 23)
:
In the 1st equation, replace b wtih (-s + 23)
(-s + 23) + m = 20
-s + m = 20 - 23
-s + m = -3
:
Add the above equation to the 2nd equation
-s + m = -3
s + m = 29
-------------- addition eliminates s, find m
2m = 26
m = 13 yrs is Mike's age
then
b + 13 = 20
b = 20 - 13
b = 7 yrs is Bill's age
and
7 + s = 23
s = 23 - 7
s = 16 is Sally's age
:
" The difference in age between the old child and young child is what?
16 - 7 = 9 yrs is the diff from oldest and youngest